Part A 0.90g of hydrogen chloride (HCl ) is dissolved in water to make 4.0L of solution. What is the pH of the resulting hydrochloric acid solution? Part B 0.75g of sodium hydroxide (NaOH ) pellets are dissolved in water to make 4.5L of solution. What is the pH of this solution?
Answer #1
molar mass of HCl is 36.5 thus moles in 0.90 g HCl = 0.0247 mole thus molarity of HCl = [H+] = 0.0247 mole / 4.0 L = 0.0062 M pH = - log [H+] = -log 0.0062 = 2.2 --------------------------------------------- molar mass of NaOH is 40.0 g/ mol thus moles in 0.75 g NaOH = 0.75 / 40.0 = 0.0188 mole [NaOH] = [OH-] = 0.0188 / 4.5 = 0.0042 M pOH = - log [OH-] = - log 0.0042 = 2.4 pH = 14 -pOH = 14 - 2.4 = 11.6
Answer #2

no. of moles of NaOH = 0.75/40 = 0.0187

so, concentration = 0.0187/4.5 = 0.00417 M

pOH = -log[OH-] = 2.38

so, pH= 14 - pOH = 11.6